OpenStudy (anonymous):

If the sin 45° is 0.707, then the cos _____° is 0.707.

OpenStudy (anonymous):

We know that sin squared plus cos squared = 1. We also know the values for sin at 45 degrees, so we can see what the cos is at that angle. Alternatively, you can see the answer clearly on a unit circle.

OpenStudy (anonymous):

Im still confused can you explain a little more. Sorry.

OpenStudy (anonymous):

If sin = 0.707 (rounded square root of 1/2), then \[\sin^2(x) = \frac{1}{2}\] Therefore, \[\sin^2 + \cos^2 = 1\] \[\frac{1}{2} + \cos^2(x) = 1\] \[\cos(x) = \frac{1}{\sqrt{2}} \approx 0.707 \]

OpenStudy (anonymous):

Therefore, since the sin(45) = 0.707, the cos(45) = 0.707

OpenStudy (anonymous):

oh ok thanks